OpenStudy (anonymous):

2x-3y=9 in slope intercept form?

pooja195 (pooja195):

2x-3y=9 The slope-intercept form is: y=mx+b So, isolate the y. Step 1: Subtract 2x from both sides. -3y=9-2x -3y=-2x+9 Step 2: Divide by (-3). y=(-2/-3)x+(9/-3) y=(2/3)x-3 Step 3: Identify the slope and y-intercept. y=mx+b m=2/3 b=-3

OpenStudy (anonymous):

Slope-intercept form is this: y = mx + b We can rearrange the equation so that it's in y = mx + b form which is the slope-intercept form. 2x - 3y = 9 - 3y = -2x + 9 <-- Here I brought 2x over to the other side. y = (-2x + 9)/-3 <-- Here I divided both sides by -3 so y is left alone. y = 2/3x - 3 <-- Here, I divided the components of the bracket by -3 to get 2/3x and -3 and we have achieved our slope-intercept form.

OpenStudy (anonymous):

thx! pooja195 and genius12

pooja195 (pooja195):

np

pooja195 (pooja195):

:)